# BEAVER: Building Environments with Assessable Variation for Evaluating Multi-Objective Reinforcement Learning

### v1

@@ Abstract
Recent years have seen significant advancements in designing reinforcement learning (RL)-based agents for building energy management. While individual success is observed in simulated or controlled environments, the scalability of RL approaches in terms of efficiency and generalization across building dynamics and operational scenarios remains an open question. In this work, we formally characterize the generalization space for the cross-environment, multi-objective building energy management task, and formulate the multi-objective contextual RL problem. Such a formulation helps understand the challenges of transferring learned policies across varied operational contexts such as climate and heat convection dynamics under multiple control objectives such as comfort level and energy consumption. We provide a principled way to parameterize such contextual information in realistic building RL environments, and construct a novel benchmark to facilitate the evaluation of generalizable RL algorithms in practical building control tasks. Our results show that existing multi-objective RL methods are capable of achieving reasonable trade-offs between conflicting objectives. However, their performance degrades under certain environment variations, underscoring the importance of incorporating dynamics-dependent contextual information into the policy learning process.
Machine Learning, ICML
@@ 1 Introduction
To ground learning-based decision-making for building systems, trained agents face similar challenges to those of the building operator: diverse objectives, varying environments (Nweye et al., 2023). Reinforcement learning (RL) algorithms hold promise in interacting with complex unknown building dynamics and implementing high-dimensional control actions for various HVAC devices. These algorithms have been shown to be effective in optimizing HVAC operations for energy efficiency or occupant comfort (Goldfeder & Sipple, 2023; Yu et al., 2021). However, two key challenges remain underexplored in this context: the ability of RL agents to generalize across diverse building dynamics, and the need to simultaneously optimize multiple, often conflicting objectives. Most existing RL-based controllers require numerous samples, and are trained and evaluated on the same building configuration, which limits their applicability in real-world scenarios where deployment environments may differ (Teoh et al., 2025; Xu et al., 2020). Furthermore, real-world building control tasks often require balancing multiple objectives, such as minimizing energy consumption while maximizing indoor comfort (Wang et al., 2021). Since each occupant or controller may have distinct preference vectors over these objectives, it is crucial to develop RL agents that are sample-efficient, generalizable, and effective across diverse reward functions and dynamics.
@@
From the RL algorithmic side, existing multi-objective RL (MORL) research largely assumes a single and static environment, while it is unclear how MORL performs across varying underlying dynamics (Liu et al., 2025; Yang et al., 2019; Rame et al., 2023). Some research discusses transfer learning and generalization in building RL (Fang et al., 2023; Deng & Chen, 2021; Zhang et al., 2020), for example by pre-training in a source building and adapting a few layers of the policy network to the target building. To advance the practical adoption of RL in building systems, we further identify two concrete research questions: i). What is the space of building generalization and how to appropriately model the space? ii). How do cross-building RL agents behave under varying reward preferences?
@@
In this work, we build upon a realistic, physics-principled RL environment for the building energy management system (Zhang et al., 2023), and design the benchmark BEAVER to support multi-objective and generalization evaluation via two important types of context variations for building systems: the variation in heat convection (which intrinsically impacts the underlying state transition) and climate zones (which directly impacts exogenous states). Our proposed method standardizes the evaluation of assessable variations in building RL tasks by automating the process of environment construction conditioned on dynamics and climate variations, reward design incorporating multiple objectives, and RL generalization evaluation. We formalize such a building RL environment as a Multi-Objective Contextual MDP (MOC-MDP), and evaluate generalization using a set of performance metrics. Preliminary evaluations on current MORL algorithms indicate their limited capabilities in addressing diverse variations of building environments. The code is available at https://github.com/RuohLiuq/BEAVER.git.
@@ 2 Multi-Objective Contextual MDP
Multi-objective contextual Markov decision process (MOC-MDP) To study generalization in MORL, we adopt the formulation of MOC-MDP as introduced by  (Teoh et al., 2025). MOC-MDP is represented by the tuple $<\mathcal{C,S,A,P,R}_{1:n},\Omega,f,\gamma,\boldsymbol{M}>$, where $\mathcal{C}$ denotes the context space, and $\boldsymbol{M}$ is a function mapping any $c\in\mathcal{C}$ to a MOC-MDP, i.e. $\boldsymbol{M}(c)=\langle\mathcal{S},\mathcal{A},\mathcal{P}^{c},\mathcal{R}_{1:n},\Omega,f,\gamma,\boldsymbol{M}\rangle$.
@@
Evaluation metrics. BEAVER supports both quantitative evaluation and visual analysis. For quantitative evaluation, BEAVER key performance metrics, including Hypervolume (HV), Sparsity (SP), and Expected Utility (EU), to assess the generalization capability of each training mode. The detailed definitions of evaluation metrics are provided in Appendix A. Briefly, a higher HV indicates a better approximation of the Pareto front, capturing both convergence and diversity. EU measures the average utility over a predefined distribution of preferences and higher values indicate better overall policy performance. In contrast, a smaller SP reflects a denser set of solutions along the Pareto front.
@@ 3 Benchmark Design: Generalization in Building Control Tasks
Building dynamics. We adopt the classical Resistance–Capacitance (RC) network to model the thermal dynamics and set up a modular RL environment of a building (Zhang et al., 2023; Ma et al., 2012). For a building of $M$ zones, the zonal temperature $T_{i}$’s dynamics can be described by the following first-order differential equation:
@@
Multi-Objective Reward Design. As indicated by the example in Introduction, there exist diverse rewards for building RL. Thermal comfort (e.g., temperature deviation from setpoints) and energy consumption are commonly used components in the reward design of many existing benchmarks and control methods for building HVAC systems (Nweye et al., 2023; Zhang et al., 2023; Goldfeder & Sipple, 2024). For the brevity of this paper, we list two examples of customized reward functions at each timestep $t$:
@@
Dynamics and Climate Variations. To model variations in building’s heat dynamics, BEAVER proposes to model the $U_{\text{wall}}$ parameter for wall’s U-factors. In the thermal modeling practice, the U-factor is a measure of a wall’s ability to insulate with units of $W/m^{2}\cdot^{\circ}C$. The lower U-factors indicate better insulation and less heat transfer. Indeed, each building has different thermal configurations affected by materials, wall thickness and , leading to varying $C_{i}$ and $R_{i}$ parameters in (2). In BEAVER, we follow the reference buildings of EnergyPlus provided by U.S. Department of Energy (Deru et al., 2011), and sample a group of practical $U_{\text{wall}}$ values, which is further converted to each wall’s $R_{i},C_{i}$ values. BEAVER also enables easy user configurations to train and evaluate on varying building dynamics (See Appendix E).
@@
Note that in our setting, the context is not observable to the agent. This is because if the context were observable beforehand, it could be concatenated with the state to form an augmented state space, reducing the MOC-MDP to a universal MOMDP (Teoh et al., 2025).
@@ 4 Experiments and Evaluation
HV$(10^{7})$uparrow 8.78$\pm$0.11 8.74$\pm$0.10 8.59$\pm$0.09 9.59$\pm$0.05 10.00$\pm$0.02 10.13$\pm$0.14
EU$(10^{3})$uparrow 8.80$\pm$0.06 8.78$\pm$0.06 8.71$\pm$0.05 9.25$\pm$0.04 9.58$\pm$0.02 9.67$\pm$0.02
SP$(10^{5})$downarrow 0.79$\pm$0.37 0.80$\pm$0.30 7.24$\pm$0.38 0.31$\pm$0.09 0.19$\pm$0.08 0.19$\pm$0.09
In this Section, we demonstrate how to evaluate generalization performance using our benchmark with state-of-the-art MORL algorithms. We adopt C-MORL (Liu et al., 2025) as our baseline method, which solves a policy optimization problem constrained by a the limits over a group of preference vectors. Studies implemented in C-MORL show strong performance in multi-objective tasks, particularly with a large number of objectives. Our benchmark is also fully compatible with the MORL-Generalization framework (Teoh et al., 2025), which integrates a collection of multi-objective environments (LavaGrid, MuJoCo and etc) with diverse contextual variations and implementations of various MORL algorithms. This allows users to easily evaluate existing methods or their own custom algorithms.
@@ 5 Conclusion and Future Directions
Appendix A Experiment Setup Details
A.1 MORL Evaluation Metrics
We evaluate the quality of Pareto front with the following metrics:
Definition A.1.
[…]
Definition A.2.
(Expected Utility). Let $P$ be a Pareto front approximation in an $n$-dimensional objective space and $\Pi$ be the policy set. The Expected Utility metric is $\mathcal{U}(P)$ $:\mathcal{U}(P)=\mathbb{E}_{\boldsymbol{\omega}\sim\Omega}\left[\boldsymbol{\omega}^{\top}\boldsymbol{G}^{\pi^{SMP}_{\boldsymbol{\omega}}}\right].$ A higher expected utility is better.
Definition A.3.
(Sparsity). Let $P$ be a Pareto front approximation in an $n$-dimensional objective space. Then the Sparsity metric $\mathcal{S}(P)$ is
$\mathcal{S}(P)=\frac{1}{|P|-1}\sum_{i=1}^{n}\sum_{k=1}^{|P|-1}\left(\tilde{G}_{i}(k)-\tilde{G}_{i}(k+1)\right)^{2},$ (3)
where $\tilde{G}_{i}$ is the sorted list for the $i^{th}$ objective values in $P$, and $\tilde{G}_{i}(k)$ is the $k^{th}$ value in this sorted list. Lower sparsity is better.
All three HV, EU, and SP metrics are integrated in MOC-MDP evaluation of BEAVER, which can be called to assess the performance of different policies in the multi-objective setting.
Appendix B Modeling of Dynamics Variations
In BEAVER, the building thermal control environment that controls the temperature of a commercial building with $M$ zones. The states contain the temperature of multiple zones, the heat acquisition from occupant’s activities, the heat gain from the solar irradiance, the ground temperature, the outdoor environment temperature, and the current electricity price $c[t]$. The actions set the controlled heating supply of the zones. The conflicting objectives can be minimizing energy cost, reducing the temperature difference between zonal temperatures and user-set points. We also set up additional reward functions in BEAVER, such as managing the ramping rate of power consumption:
$\mathcal{R}_{ramp}=M-|(\sum_{i}|P_{i}[t]|-\sum_{i}|P_{i}[t-1]|)|.$
To model the dynamics variations, we use EnergyPlus as a reference for standard building simulations. To identify the U-wall values, we simulate all reference buildings and find the these distinct $U_{wall}$ values in “Envelope Summary” table in the main HTML output file generated by EnergyPlus. This table provides a comprehensive overview of the thermal properties of the building envelope, including the U-values for all constructed elements such as walls, roofs, windows, and floors. We make use of EnergyPlus reference buildings to extract $U_{wall}$. They are developed by the U.S. Department of Energy, are a comprehensive set of standardized building models used for energy analysis and simulation. These models represent various commercial building types across different climate zones and are accompanied by extensive documentation and the necessary input files for running simulations in EnergyPlus.
Based on the $U_{\text{wall}}$ extracted from EnergyPlus reference buildings, we define a sampling range for each individual thermal parameter in $U_{\text{wall}}$. During environment construction, we independently sample these U-values within their respective ranges to simulate various building envelopes. While the rest of the building configuration (e.g., layout, HVAC settings) remains consistent with a chosen reference building, this strategy allows us to approximate realistic and cross-building variations. In future work, we plan to expand this sampling approach to additional parameters such as spatial layout and occupancy.
Appendix C MORL Algorithms
We choose to implement C-MORL algorithm due to its strong performance and scalability to larger number of objectives (Liu et al., 2025). In the C-MORL implementation, during the policy initialization stage, we employ parallel training of $M$ policies for pre-known preference vector $\mathbf{\omega}$. We maintain a policy buffer, meaning that in addition to the final policy obtained in the Pareto initialization intermediate policies are also saved in the buffer. We conduct policy selection in both after the Pareto initialization stage and during the Pareto extension stage. Then in the Pareto extension stage, we achieve the goal of filling the Pareto front from selected solutions $\mathcal{X}_{extension}$ toward different directions by solving constrained optimization on the selected policies. Specifically, we consider solving the following problem in which return constraints are controlled by a hyperparameter $\beta\in(0,1)$:
$\pi_{r+1,i}=\arg\max_{\pi\in\Pi_{\theta}}\{G_{l}^{\pi}:G_{i}^{\pi}\geq\beta G_{i}^{\pi_{r}},i=1,\ldots,n,i\neq l\},$ (4)
We note that most of MORL algorithms to date focus on static environment, where training and evaluation are implemented in a single environment with fixed state transitions or data distributions. While research in multi-task RL and representation learning in RL have proposed to use contextual MDP (CMDP) to model policy transfer, multi-task learning, and environment generalization, previous works focus on single reward preference (Sodhani et al., 2021; Teh et al., 2017). This work intends to identify the notable gap in the literature regarding the simultaneous consideration of multi-objectivity and generalization across contexts (i.e. environments or tasks) in building RL tasks.
Appendix D Discussion of Limitations
Restrictions on State and Action Space. In our current design of BEAVER, we train and evaluate RL policies with the same dimension of state and action spaces. This design may affect the generalization capabilities of the trained policies, as real-world buildings can have different number of zones and often involve varying but finite dimensions, which can create a mismatch between learned and actual environments. Future work could explore adaptive mechanisms to handle such variability, such as dimensionality reduction techniques and state- or action-level feature embeddings (Bitzer et al., 2010; Chandak et al., 2019).
Construction of Building Dynamical Models. Discovering and building up thermal dynamics buildings for building systems is nontrivial, and previous works have focused on model-based approaches to set up high-fidelity model (Crawley et al., 2001; Wetter et al., 2014). In the current design of BEAVER, we mainly treat climate and $R,C$ values as two major sources of variations in building models, representing exogenous and dynamical factors respectively. And each building is unique, due to its physical layout, equipment configurations, and location. It is interesting to further customize high fidelity simulation to a specific target building considering detailed building structure, materials, and geographical factors.
Appendix E BEAVER Additional Implementation Examples
In this section, we provide a brief introduction to key implementation components of the BEAVER benchmark. The following code snippets illustrate how to instantiate environments, configure dynamic parameters, define climate profiles, and implement vectorized reward functions.
The code below shows how to create a BEAVER environment by generating building parameters and passing them to the simulation environment. Users can flexibly configure simulation conditions, including building type, climate profiles, and locations, through the ‘ParameterGenerator‘ interface.
1params = ParameterGenerator(Building=args.building, Weather=weather, Location=args.location)
2env = BuildingEnv_DR_2d(params)
Listing 1: Creating a BEAVER environment with building and weather specifications
This snippet defines the dynamic building components and their corresponding ranges for dynamic parameter sampling (e.g., $U_{Wall}$ values).
1# Index-to-name mapping for U_Wall elements
2self.dyn_ind_to_name = {
3 0: ’intwall’,
4 1: ’floor’,
5 2: ’outwall’,
6 3: ’roof’,
7 4: ’ceiling’,
8 5: ’groundfloor’,
9 6: ’window’
10 }
11
12bounds_range = {
13 ’intwall’: (0.774, 6.299),
14 ’floor’: (0.386, 3.145),
15 ’outwall’: (0.269, 2.191),
16 ’roof’: (0.160, 1.304),
17 ’ceiling’: (0.386, 3.145),
18 ’groundfloor’: (0.386, 3.145),
19 ’window’: (1.950, 3.622)
20 }
Listing 2: Sample code for specifying dynamic parameters
BEAVER includes a set of built-in climate files, and users can specify desired weather profiles to evaluate the generalization performance of their learned policies under diverse climatic conditions.
1weather_list = [’Mixed_Marine’, ’Cool_Marine’, ’Warm_Humid’, ’Warm_Dry’, ’Hot_Humid’]
Listing 3: Sample code for specifying weather profile
This example shows how BEAVER environments return multi-objective rewards, including user comfort and saving energy cost, which are computed from indoor temperature difference and energy cost. Users can define custom multi-objective reward functions or select from the predefined options provided by BEAVER.
1def default_reward_function(self, state, action, error, state_new, step_idx):
2 reward_comfort = (20*self.roomnum-LA.norm(error, 1))/20
3 reward_price = self.roomnum-self.Price_factor*self.ElectricityPrice[step_idx]*LA.norm(action, 1)
4 return reward_comfort, reward_price
Listing 4: Sample code for defining vectorized reward

### v3

@@ Abstract
Recent years have seen significant advancements in designing reinforcement learning (RL)-based agents for building energy management. While individual success is observed in simulated or controlled environments, the scalability of RL approaches in terms of efficiency and generalization across building dynamics and operational scenarios remains an open question. In this work, we formally characterize the generalization space for the cross-environment, multi-objective building energy management task, and formulate the multi-objective contextual RL problem. Such a formulation helps understand the challenges of transferring learned policies across varied operational contexts such as climate and heat convection dynamics under multiple control objectives such as comfort level and energy consumption. We provide a principled way to parameterize such contextual information in realistic building RL environments, and construct a novel benchmark to facilitate the evaluation of generalizable RL algorithms in practical building control tasks. Our results show that existing multi-objective RL methods are capable of achieving reasonable trade-offs between conflicting objectives. However, their performance degrades under certain environment variations, underscoring the importance of incorporating dynamics-dependent contextual information into the policy learning process. BEAVER is open sourced for building RL developments  [^1] https://github.com/chennnnnyize/BEAVER.
[…]
Yize Chen
@@ 1 Introduction
To ground learning-based decision-making for building systems, trained agents face similar challenges to those of the building operator: diverse objectives, varying environments Nweye et al. (2023). Reinforcement learning (RL) algorithms hold promise in interacting with complex unknown building dynamics and implementing high-dimensional control actions for various HVAC devices. These algorithms have been shown to be effective in optimizing HVAC operations for energy efficiency or occupant comfort Goldfeder & Sipple (2023); Yu et al. (2021). However, two key challenges remain underexplored in this context: the ability of RL agents to generalize across diverse building dynamics, and the need to simultaneously optimize multiple, often conflicting objectives. Most existing RL-based controllers require numerous samples, and are trained and evaluated on the same building configuration, which limits their applicability in real-world scenarios where deployment environments may differ Teoh et al. (2025); Xu et al. (2020). Furthermore, real-world building control tasks often require balancing multiple objectives, such as minimizing energy consumption while maximizing indoor comfort Wang et al. (2021). Since each occupant or controller may have distinct preference vectors over these objectives, it is crucial to develop RL agents that are sample-efficient, generalizable, and effective across diverse reward functions and dynamics.
@@
From the RL algorithmic side, existing multi-objective RL (MORL) research largely assumes a single and static environment, while it is unclear how MORL performs across varying underlying dynamics Liu et al. (2025); Yang et al. (2019); Rame et al. (2023). Some research discusses transfer learning and generalization in building RL Fang et al. (2023); Deng & Chen (2021); Zhang et al. (2020), for example by pre-training in a source building and adapting a few layers of the policy network to the target building. To advance the practical adoption of RL in building systems, we further identify two concrete research questions: i). What is the space of building generalization and how to appropriately model the space? ii). How do cross-building RL agents behave under varying reward preferences?
@@
In this work, we build upon a realistic, physics-principled RL environment for the building energy management system Zhang et al. (2023), and design the benchmark BEAVER to support multi-objective and generalization evaluation via two important types of context variations for building systems: the variation in heat convection (which intrinsically impacts the underlying state transition) and climate zones (which directly impacts exogenous states). Our proposed method standardizes the evaluation of assessable variations in building RL tasks by automating the process of environment construction conditioned on dynamics and climate variations, reward design incorporating multiple objectives, and RL generalization evaluation. We formalize such a building RL environment as a Multi-Objective Contextual MDP (MOC-MDP), and evaluate generalization using a set of performance metrics. Preliminary evaluations on current MORL algorithms indicate their limited capabilities in addressing diverse variations of building environments. The code is available at https://github.com/chennnnnyize/BEAVER.git.
@@ 2 Multi-Objective Contextual MDP
Multi-objective contextual Markov decision process (MOC-MDP) To study generalization in MORL, we adopt the formulation of MOC-MDP as introduced by  Teoh et al. (2025). MOC-MDP is represented by the tuple $<\mathcal{C,S,A,P,R}_{1:n},\Omega,f,\gamma,\boldsymbol{M}>$, where $\mathcal{C}$ denotes the context space, and $\boldsymbol{M}$ is a function mapping any $c\in\mathcal{C}$ to a MOC-MDP, i.e. $\boldsymbol{M}(c)=\langle\mathcal{S},\mathcal{A},\mathcal{P}^{c},\mathcal{R}_{1:n},\Omega,f,\gamma,\boldsymbol{M}\rangle$.
@@
Evaluation metrics. BEAVER supports both quantitative evaluation and visual analysis. For quantitative evaluation, BEAVER key performance metrics, including Hypervolume (HV), Sparsity (SP), and Expected Utility (EU), to assess the generalization capability of each training mode. The detailed definitions of evaluation metrics are provided in Appendix 6. Briefly, a higher HV indicates a better approximation of the Pareto front, capturing both convergence and diversity. EU measures the average utility over a predefined distribution of preferences and higher values indicate better overall policy performance. In contrast, a smaller SP reflects a denser set of solutions along the Pareto front.
@@ 3 Benchmark Design: Generalization in Building Control Tasks
Building dynamics. We adopt the classical Resistance–Capacitance (RC) network to model the thermal dynamics and set up a modular RL environment of a building Zhang et al. (2023); Ma et al. (2012). For a building of $M$ zones, the zonal temperature $T_{i}$’s dynamics can be described by the following first-order differential equation:
@@
Multi-Objective Reward Design. As indicated by the example in Introduction, there exist diverse rewards for building RL. Thermal comfort (e.g., temperature deviation from setpoints) and energy consumption are commonly used components in the reward design of many existing benchmarks and control methods for building HVAC systems Nweye et al. (2023); Zhang et al. (2023); Goldfeder & Sipple (2024). For the brevity of this paper, we list two examples of customized reward functions at each timestep $t$:
@@
Dynamics and Climate Variations. To model variations in building’s heat dynamics, BEAVER proposes to model the $U_{\text{wall}}$ parameter for wall’s U-factors. In the thermal modeling practice, the U-factor is a measure of a wall’s ability to insulate with units of $W/m^{2}\cdot^{\circ}C$. The lower U-factors indicate better insulation and less heat transfer. Indeed, each building has different thermal configurations affected by materials, wall thickness and , leading to varying $C_{i}$ and $R_{i}$ parameters in (2). In BEAVER, we follow the reference buildings of EnergyPlus provided by U.S. Department of Energy Deru et al. (2011), and sample a group of practical $U_{\text{wall}}$ values, which is further converted to each wall’s $R_{i},C_{i}$ values. BEAVER also enables easy user configurations to train and evaluate on varying building dynamics (See Appendix 10).
@@
Note that in our setting, the context is not observable to the agent. This is because if the context were observable beforehand, it could be concatenated with the state to form an augmented state space, reducing the MOC-MDP to a universal MOMDP Teoh et al. (2025).
@@ 4 Experiments and Evaluation
[…]
In this Section, we demonstrate how to evaluate generalization performance using our benchmark with state-of-the-art MORL algorithms. We adopt C-MORL Liu et al. (2025) as our baseline method, which solves a policy optimization problem constrained by a the limits over a group of preference vectors. Studies implemented in C-MORL show strong performance in multi-objective tasks, particularly with a large number of objectives. Our benchmark is also fully compatible with the MORL-Generalization framework Teoh et al. (2025), which integrates a collection of multi-objective environments (LavaGrid, MuJoCo and etc) with diverse contextual variations and implementations of various MORL algorithms. This allows users to easily evaluate existing methods or their own custom algorithms.